# Robust model predictive control exploiting monotonicity properties

### v2

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ IV-A Generalization via mixed-monotonicity
$\exists{u}_{\textnormal{RCIS}}^{\left[1:\mu_{s}\right]}\in{\mathbb{U}}:{d}({x}_{\textnormal{RCIS}}^{s\pm},{x}_{\textnormal{RCIS}}^{s\mp},{u}_{\textnormal{RCIS}}^{s},{p}^{\pm},{p}^{\mp})\in{\mathbb{X}}_{f},\ \forall s\in{\mathbb{S}}.$
@@ V Nonlinear, non-monotone CSTR Cascade
Educt $A$ reacts with $B$ to form $R$ with the second order side reaction of $2A\rightarrow S$. The ODEs of the $i$th CSTR are with the reaction rates $r_{1,i}=k_{1,i}\exp{-\frac{E_{A1,i}}{R_{\text{gas}}(T_{R,i}+273.15)}}c_{A,i}c_{B,i}$ and $r_{2,i}=k_{2,i}\exp{-\frac{E_{A2,i}}{R_{\text{gas}}(T_{R,i}+273.15)}}c_{A,i}^{2}$
@@
$\ell({x}_{k},{u}_{k},{u}_{k-1})=({x}_{k}-{x}_{\text{set}})^{\intercal}{Q}({x}_{k}-{x}_{\text{set}})\\ +({u}_{k}-{u}_{\text{set}})^{\intercal}{R}({u}_{k}-{u}_{\text{set}})+({u}_{k}-{u}_{k-1})^{\intercal}{R}_{d}({u}_{k}-{u}_{k-1})\\ +(\sum u_{A,i,k}-1.5)^{2}++(\sum u_{B,i,k}-1.5)^{2}.$ (11)

### v3

@@ Abstract
This work has been received on 24 January 2025 and has been accepted by the IEEE for publication with the Digital Object Identifier: 10.1109/TAC.2025.3558137 on 29 March 2025. ©2025 The Authors. This work is licensed under a Creative Commons Attribution 4.0 License. For more information, see http://creativecommons.org/licenses/by/4.0/
@@ IV-A Generalization via mixed-monotonicity
$\exists{u}_{\textnormal{RCIS}}^{\left[1:\mu_{s}\right]}\in{\mathbb{U}}:{d}({x}_{\textnormal{RCIS}}^{s\pm},{p}^{\pm},{u}_{\textnormal{RCIS}}^{s},{x}_{\textnormal{RCIS}}^{s\mp},{p}^{\mp})\in{\mathbb{X}}_{f},\ \forall s\in{\mathbb{S}}.$
@@ V Nonlinear, non-monotone CSTR Cascade
Educt $A$ reacts with $B$ to form $R$ with the second order side reaction of $2A\rightarrow S$. The differential equations of the $i$th CSTR are with the reaction rates $r_{1,i}=k_{1,i}\exp{-\frac{E_{A1,i}}{R_{\text{gas}}(T_{R,i}+273.15)}}c_{A,i}c_{B,i}$ and $r_{2,i}=k_{2,i}\exp{-\frac{E_{A2,i}}{R_{\text{gas}}(T_{R,i}+273.15)}}c_{A,i}^{2}$
@@
$\ell({x}_{k},{u}_{k},{u}_{k-1})=({x}_{k}-{x}_{\text{set}})^{\intercal}{Q}({x}_{k}-{x}_{\text{set}})\\ +({u}_{k}-{u}_{\text{set}})^{\intercal}{R}({u}_{k}-{u}_{\text{set}})+({u}_{k}-{u}_{k-1})^{\intercal}{R}_{d}({u}_{k}-{u}_{k-1})\\ +(\sum u_{A,i,k}-1.5)^{2}+(\sum u_{B,i,k}-1.5)^{2}.$ (10)